OpenStudy (ffrocks3):

What is the solution set for |-4x - 4| - 28 = -4? a. x = -7 and x = 5 b. x = -7 and x = -5 c. x = 7 and x = 5 d. x = 7 and x = -5

OpenStudy (solomonzelman):

\(\large\color{black}{ \left| -4x-4\right| -28=-4 }\) you are allowed to factor 4 out of the absolute value (but don't factor -4, because \(\large\color{red}{ \left| -a\right|~ \ne ~-~\left| a\right| }\) ) \(\large\color{black}{ 4\left| -x-1\right| -28=-4 }\) divide everything in the equation by 4, \(\large\color{black}{ \left| -x-1\right| -7=-1 }\) add 7 to both sides \(\large\color{black}{ \left| -x-1\right| -7\color{blue}{+7}=-1\color{blue}{+7} }\)

OpenStudy (solomonzelman):

can you take it from here, or need more help?

OpenStudy (ffrocks3):

I think I got it thanks :)

OpenStudy (solomonzelman):

when you are done, just tell me the answer you get, to make sure you have it correctly, okay?

OpenStudy (ffrocks3):

ok

OpenStudy (ffrocks3):

I need more help sorry.

OpenStudy (ffrocks3):

I did this \[|-x-1|=6 \] Would it then just be: x=5 or x=7? idrk

OpenStudy (solomonzelman):

Look, this is what you are supposed to do, \(\large\color{black}{ \left| -x-1=6 \right| }\) Gives that, \(\large\color{black}{ -x-1=6 }\) or \(\large\color{black}{ -x-1=-6 }\) \(\large\color{black}{ -x-1\color{red}{+1}=6 \color{red}{+1}}\) \(\large\color{black}{ -x-1\color{red}{+1}=-6 \color{red}{+1}}\) \(\large\color{black}{ -x=5}\) \(\large\color{black}{ -x=-5}\)

OpenStudy (solomonzelman):

and the last steps you already know.

OpenStudy (ffrocks3):

oh ok Thank You for the help!

OpenStudy (solomonzelman):

anytime:)